Solve for b. a = 1/2 bc^2
This is really strange and I am not sure what to do? Where are all the numbers?

Answer by stanbon(75887)

Solve for b. a = 1/2 bc^2
-----------
Comment: When you are studying algebra letters ARE numbers.
----------
Your Problem:
a = (1/2)bc^2
Multiply both sides by 2 to get:
2a = bc^2
Divide both sides by c^2 to solve for "b":
b = (2a)/b^2
=============
cheers,
Stan H.